# VoxEval: Benchmarking the Knowledge Understanding Capabilities of End-to-End Spoken Language Models

### v3

@@ Appendix A Explanation for Various Speech Recording Conditions
Appendix B Additional Experimental Settings
For models used in our evaluation, SpeechGPT, TWIST, SPIRIT-LM, and Moshi contain 7B parameters, and GLM-4-Voice contains 9B parameters. We use the base version of SPIRIT-LM. All experiments are conducted on a single NVIDIA RTX A6000 GPU. For all the models, we use the default hyperparameters provided in their code implementations. Each experiment is performed once.
Appendix C Case Study for SLM Responses
Table 10 presents example SLM responses to questions from VoxEval as well as our comments.
@@ Appendix C Case Study for SLM Responses
Appendix D String Matching vs. LLM Extraction
@@
3:  for each input audio condition setting $S_{i}$ do
5:     Initialize an empty list $D_{i}=[]$ to store difference scores for $S_{i}$.
6:     for each subject $n\in\{1,2,\dots,N\}$ do
8:        Compute the maximum absolute score difference:
9:        Append $d_{n}$ to $D_{i}$: $D_{i}\leftarrow D_{i}\cup\{d_{n}\}$.
10:     end for
11:     Use $D_{i}=\{d_{1},d_{2},\dots,d_{56}\}$ to generate a box plot for $S_{i}$.
12:  end for

### v4

@@ Appendix E Explanation for Various Speech Recording Conditions
Noise Type Settings
Gaussian Noise min_snr_db=5.0
max_snr_db=20.0
Colored Noise min_snr_db=5.0
max_snr_db=20.0
Background Music min_snr_db=3.0
max_snr_db=10.0
Short Interruption Noise min_snr_db=3.0
max_snr_db=10.0
Table 13: Transformation settings for the various types of noise used.
@@
Appendix F Additional Experimental Settings
For models used in our evaluation, SpeechGPT, TWIST, SPIRIT-LM, and Moshi contain 7B parameters, and GLM-4-Voice contains 9B parameters. We use the base version of SPIRIT-LM. For all the models, we use the default hyperparameters provided in their code implementations. Each experiment is performed once.
Appendix G Case Study for SLM Responses
Table 15 presents example SLM responses to questions from VoxEval as well as our comments. It highlights the most characteristic answer style for each SLM. We randomly choose responses from each SLM under the Alloy setting while ensuring that the selected answer style is the most representative of that particular SLM.
@@ Appendix G Case Study for SLM Responses
Appendix H String Matching vs. LLM Extraction